# Keys to Robust Edits: From Theoretical Insights to Practical Advances

### v1

@@ Abstract
Large language models (LLMs) have revolutionized knowledge storage and retrieval, but face challenges with conflicting and outdated information. Knowledge editing techniques have been proposed to address these issues, yet they struggle with robustness tests involving long contexts, paraphrased subjects, and continuous edits. This work investigates the cause of these failures in locate-and-edit methods, offering theoretical insights into their key-value modeling and deriving mathematical bounds for robust and specific edits, leading to a novel ’group discussion’ conceptual model for locate-and-edit methods. Empirical analysis reveals that keys used by current methods fail to meet robustness and specificity requirements. To address this, we propose a Robust Edit Pathway (REP) that disentangles editing keys from LLMs’ inner representations. Evaluations on LLaMA2-7B and Mistral-7B using the CounterFact dataset show that REP significantly improves robustness across various metrics, both in-domain and out-of-domain, with minimal trade-offs in success rate and locality. Our findings advance the development of reliable and flexible knowledge updating in LLMs.
@@ 1 Introduction
Large language models (LLMs, OpenAI 2023; Touvron et al. 2023a; b) have revolutionized the storage and retrieval of vast amounts of human knowledge. However, their training on extensive internet-based datasets inevitably leads to the incorporation of conflicting and rapidly outdated information, potentially contributing to the phenomenon of model hallucination (Zhang et al., 2023). To mitigate such issues, researchers have proposed knowledge editing techniques (Sinitsin et al., 2020; De Cao et al., 2021), which employ carefully designed, constrained updates to specific parameters. While giving promising results (Mitchell et al., 2022a; Meng et al., 2023a), there still exist challenges in its implementation, particularly their susceptibility to failure with robustness tests, e.g., when faced with long contexts, paraphrased subjects, and continuous edits (Ma et al., 2024c; Yang et al., 2024). For example, in Figure 1, although one can edit the knowledge ‘Slovenia belongs to the continent of’ from ‘Europe’ to ‘Antarctica’, the model still falls into failure when the subject is rephrased to ‘The Republic of Slovenia’.
We aim to investigate the cause of such editing failures and further resolve them, with a specific focus on the widely recognized locate-and-edit methods (Meng et al., 2023a; Li et al., 2024a; Hu et al., 2024b; Meng et al., 2023b), which use causal interventions to pinpoint factual knowledge in the early-middle MLP layers of LLMs, treating these modules as associative key-value memories (Kohonen, 1972) and constructing new memories for editing. We give a new theoretical perspective into key-value modeling, demonstrating that any value retrieved from the module can be viewed as a linear combination of existing memories, where the weights are determined by similarities of the keys. As a result, knowledge insertion can be viewed as a localized patch rather than a comprehensive rewrite. We further provide a novel conceptual model, ‘group discussion’. In this model, we view knowledge editing as inserting influence of a new voice among a mixture of opinion holders, and effective knowledge insertion requires a strong voice (i.e., large value) at the correct position (i.e., similar keys) while balancing the risk of affecting nearby group (i.e., related knowledge).
@@
In light of the group discussion view, we empirically analyze the keys used by locate-and-edit methods and find they cannot follow the ideal scenario. In particular, we find that keys of robustness queries such as the same subject (e.g., Slovenia) with different surface forms (e.g., The Republic of Slovenia), shuffled token ordering (e.g., ia Sloven) and appending long contexts (e.g., Tiffany … Slovenia), are highly dissimilar to the edit key in the representation space. In contrast, keys from different subjects (e.g., Slovenia vs Croatia) could be highly similar, when subjects demonstrate semantic similarities. These make it impossible for the edited knowledge to both affect robustness queries while not affecting those semantically similar but different subjects. To mitigate such problems, we propose to build a Robust Edit Pathway (REP) which disentangles the keys used by editing methods from the inner representations of LLMs. Specifically, we train an adaptor that applies a gate mechanism to separate the keys from the model’s inner representations, aggregating the keys for the parts that need editing, while keeping other representations intact. This adaptor is designed by aggregating semantically identical keys, and allows robust edit.
We evaluate REP on widely used LLaMA2-7B (Touvron et al., 2023b) and Mistral-7B (Jiang et al., 2023) using the CounterFact dataset, comparing it against an representative method ROME (Meng et al., 2023a). The Robust Edit Pathway leads to substantial improvements in performance, across various robustness metrics, both for in-domain and out-of-domain tests, with only a minimal trade-off in success rate and locality. These results not only validate our theoretical and empirical findings, but also demonstrate the adaptor’s ability to generalize across different types of perturbations, effectively addressing the critical robustness issues in existing knowledge editing techniques and moving closer to our goal of reliable and flexible knowledge updating in large language models. [^1] Codes can be found at https://github.com/ElliottYan/keys_to_robust_edit/
@@ 2 Related Work
As large language models have grown in complexity and size, post-modification has become increasingly challenging due to their opaque mechanisms and vast parameter spaces (Mitchell et al., 2022c; Zhong et al., 2023). This has led to heightened interest in knowledge editing, a technique for precise model modification. Knowledge editing are applied to various scenarios, such as editing for safety (Wang et al., 2024c), debias (Yan et al., 2024) and concepts (Wang et al., 2024d).
Knowledge editing approaches can be broadly categorized into two groups: those that preserve model parameters and those that modify them (Yao et al., 2023). Parameter-preserving methods include memory-based strategies (Mitchell et al., 2022d; Zhong et al., 2023; Hartvigsen et al., 2023) and alternative approaches (Dong et al., 2022; Huang et al., 2023) incorporate additional parameters into the model. On a different line of research, some work focuses on directly updating model parameters to avoid cost of an external memory, which can be further divided into meta-learning methods (Cao et al., 2021; Mitchell et al., 2022b; Tan et al., 2024) and locate-and-edit methods.
@@
Our work is in line with the locate-then-edit methods, which draw much attention as they potentially unveil how the knowledge are stored in an LLM. These approaches first identify relevant parameters before updating them to modify specific knowledge, including KnowledgeNeuron’s attribution-based neuron updating (Dai et al., 2022), ROME’s causal mediation analysis for MLP editing (Meng et al., 2023a), MEMIT’s multi-layer residual distribution (Meng et al., 2023b), PMET’s refined allocation strategy (Li et al., 2024a), and WilKE’s dynamic layer selection (Hu et al., 2024b) to reduce potential negative effects. These methods all utlize inner representations as keys for key-value modeling. In contrast, we show that inner representations cannot meet the requirements of robust and specific edits, and we propose REP to mitigate this.
Issues of Knowledge Editing.
@@
Despite the promise of knowledge editing for refining large language models, various challenges persist in practical applications. Edits often degrade general language abilities (Gu et al., 2024; Ma et al., 2024b), damage the hidden space (Wang et al., 2024b), struggle to propagate to related facts (Hua et al., 2024), and are easily forgotten during sequential updates (Gupta et al., 2024). Moreover, multi-hop reasoning can elicit old knowledge (Zhang et al., 2024), and models may collapse after few edits (Yang et al., 2024; Brown et al., 2023).
@@
Further complications include cross-lingual inconsistencies (Wang et al., 2024a), knowledge conflicts (Li et al., 2024b), and inadequate evaluation in realistic settings such as long-form generation (Rosati et al., 2024) and neighborhood knowledge (Ma et al., 2024a). These issues underscore the need for more sophisticated and comprehensive editing techniques. However, previous research largely remain focused on the outcomes of knowledge editing in various scenarios, lacking a deeper understanding of the underlying mechanisms of these algorithms and the true reasons behind their frequent failures. Our work presents both theoretical and empirical understanding regarding the reason of robustness failures of locate-and-edit methods, and proposes REP to enhance them.
Figure 2: Top: Illustration of the knowledge editing task. Middle: For a specific layer, the locate-and-edit methods use representations inside FFN module as keys and values. Bottom: The keys are diverse and dissimilar, our REP aggregates the keys to facilitate robust editing.
@@ 3 Knowledge Editing
In this section, we first formulate the task of knowledge editing and review the locate-then-edit methods.
@@
Given a knowledge triple: $(h,r,t)$, e.g., (USA, president of, Biden), knowledge editing aims to update the original target of the triple, i.e., ‘Biden’, to a new target $t_{*}$, say ‘Trump’. Recent knowledge editing methods formulate such a task with a conditional generation. Following Meng et al. (2023a) and Meng et al. (2023b), we define a knowledge $\mathbf{f}$ as a triple $(h,r,t)$ where $h$ denotes the head entity, $r$ the relation, and $t$ the tail entity (e.g., ($h=$USA, $r=$president of, $t=$Biden)).
Definition 3.1 (Task Definition of Knowledge Editing).
Given a knowledge triple $\mathbf{f}=(h,r,t)$, knowledge editing algorithm $\mathcal{A}$ aims to update the original knowledge in the language model $\mathcal{M}$ to new knowledge $\mathbf{f}^{\prime}=(h,r,t_{*})$. This task can be formally expressed as follows:
$\mathcal{M}^{\prime}=\mathcal{A}(\mathcal{M})\,,\quad\text{such that}\quad k^{\prime}\in\mathcal{M}^{\prime}\,,k\notin\mathcal{M}^{\prime}\,,$
Autoregressive transformer-based large language models (LLMs) can answer natural-language queries by leveraging implicit knowledge encoded within their parameters. For instance, given the prompt "The president of USA is", a well-trained model would likely respond with "Biden".
Building on this, the Locate-then-Edit approach, specifically in the context of ROME, demonstrates that modifying specific MLP layers locating in the critical path of an LLM suffices to edit its factual associations. This method views the MLP layers $W$ in the transformer as a form of linear associative memory, observing that any linear operation $W$ can function as a key-value store, where a set of vector keys $K=[k_{1}|k_{2}|...]$ corresponds to a set of vector values $V=[v_{1}|v_{2}|...]$.
@@
$\text{minimize}\ ||\hat{W}K-V||\ \text{such that}\ \hat{W}k_{*}=v_{*}\,,\text{by setting}\ \hat{W}=W+\boldsymbol{\Lambda}(C^{-1}k_{*})^{T}$
@@
To implement this solution, it is necessary to extract the key $k_{*}$ and calculate the value $v_{*}$.
@@
$v_{*}=\arg\min_{\mathbf{z}}\frac{1}{N}\sum_{j=1}^{N}\underbrace{-\log\mathbb{P}_{\mathcal{M}(m_{i}^{(l*)}:\mathbf{z})}[h^{\prime}|x_{j}+p]}_{\text{(a) Maximizing }h^{\prime}\text{ probability}}+\underbrace{D_{\text{KL}}\left(\mathbb{P}_{\mathcal{M}(m_{i}^{(l*)}:\mathbf{z})}[\cdot|p^{\prime}]||\mathbb{P}_{\mathcal{M}}[\cdot|p^{\prime}]\right)}_{\text{(b) Controlling essence drift}}.$
where $p^{\prime}$ is ’subject is a’.
@@ 4 Theoretical Results of Key-Value Associative Memory
The idea of keys and values in associative memory is analogous to the key-value databases in modern computer systems. What makes the difference here is that our FFNs implement a fuzzy retrieval mechanism, whereas the modern key-value databases generally require the keys to be unique.
Given a set of keys $K=[k_{1}|k_{2}|\cdots]$ and values $V=[v_{1}|v_{2}|\cdots]$, recall that locate-and-edit methods such as ROME, MEMIT, and PMET view the down project feed-forward $W$ as a linear associative memory, where it satisfies $WK\approx V$.
@@
Given $K\in\mathbb{R}^{D_{1}\times N}$ and $V\in\mathbb{R}^{D_{2}\times N}$ that are already stored in the feed-forward layer $W\in\mathbb{R}^{D_{2}\times D_{1}}$, assume $N\gg D_{1}$ and $K$ has the rank of $D_{1}$. When a new query $\hat{k}$ comes, its corresponding value can be represented as the weighted sum of existing values, $\hat{v}=\sum_{i}^{N}\alpha_{i}v_{i}$ and $\alpha=K^{T}(KK^{T})^{-1}\hat{k}$ can be solved by the Moore-Penrose pseudoinverse.
The above lemma demonstrates that the retrieved memory of a new test query can be considered as the linear combination of previously stored memory, which leads to a direct corollary.
@@
Recall that in locate-and-edit algorithms, a newly edited knowledge is injected into the memory as a new key-value pair $(k_{*},v_{*})$. Suppose the injection is lossless [^2] In a real-world scenario, the edit cannot be lossless. Here, for a clear intuition, the above lemma is presented in an ideal way as the editing process will change the value of previously stored key-value pairs. We show that even considering the lossless scenario, the current LLMs cannot satisfy robustness and specificity requirements. and we have a set of key-value pairs that represent the same semantic meaning $k_{i}\in K_{s},v_{i}\in V_{s}$ that are already stored in the memory. Then, if $K_{s}$ has full row rank, a query of $k\in K_{s}$ would retrieve a value $v=\sum_{i}\alpha_{i}v_{i}+\alpha_{*}v_{*}$.
@@
This corollary demonstrates that the editing process does not essentially ‘change’ the previously stored knowledge. Instead, it leaves the previously stored knowledge intact and counters them with a newly added value $v_{*}$.
@@
Lemma 4.4 suggests that an edited value should be first similar to the vector pointing from $t$ to $t_{*}$ after a projection with $W_{V}$. Then, the edited values $\Delta v$ should be sufficiently large to ensure that the success of the edit.
Our assumption here simplifies the connection between the edited layer and the prediction layer, as in real-world scenarios, the edit layer might pass through subsequent layers and undergo multiple attention operations before finally connecting to the prediction layer. However, the path we’re considering - from the edit layer to the prediction layer via an attention - is arguably the most direct route. We contend that this direct path is crucial and warrants particular attention and this simplification allows us to focus on the most immediate and potentially significant impact of edits.
@@
If the newly added knowledge triplet $(h,r,t_{*})$ can be robustly retrieved for all $k_{s}\in K_{s}$, it is required the following inequality to be satisfied:
$(w_{t_{*}}-w_{t})^{T}W_{V}(k_{s}^{T}C^{-1}k_{*})\cdot v_{*}^{T}>\epsilon_{1}+\epsilon_{2},\forall k_{s}\in K_{s}.$ (3)
When we look into the lemma, $\alpha_{s,*}=k_{s}^{T}C^{-1}k_{*}$ can be seen as a similarity measure on a projected space. This lemma implies that (1) $v_{*}$ is decided by $\min_{k_{s^{\prime}}\in K_{s}}(k_{s^{\prime}}C^{-1}k_{*}$), that is, $k_{*}$ should be near all $k_{s}\in K_{s}$. If not, $v_{*}$ needs to be of large magnitude to counter the difference. (2) $v_{*}$ needs not only to be aligned with the direction $(w_{t_{*}}-w_{t})$, but also has a sufficiently large magnitude to ensure prediction success.
@@
If the newly added knowledge triplet $(h,r,t_{*})$ would not be retrieved for any $k_{n}\notin K_{s}$, it is required the following inequality to be satisfied:
$(w_{n}-w_{t_{*}})^{T}W_{V}(k_{n}^{T}C^{-1}k_{*})\cdot v_{*}^{T}<\epsilon_{3},\forall k_{o}\notin K_{s}~{}and~{}\forall w\in W,$ (4)
where $t_{n}$ is the original target retrieved by $k_{n}$ and $\epsilon_{3}$ denotes the logit difference between $t_{n}$ and $t_{*}$.
One simple solution for this lemma is $k^{T}C^{-1}k_{*}=0$, which describes no superposition, as discussed in one of the concurrent work (Hu et al., 2024a). However, as superposition generally exists among existing LLMs, we discuss more general cases here. The detailed proof of all above lemmas can be found in Appendix A.
Group Discussion.
The aforementioned lemma can be understood through a simple conceptual model. Imagine each piece of knowledge as a group of individuals engaged in a discussion. Each person has an existing key-value pair in the model, where the key represents their position and the value represents the volume of their voice. When attempting to add new knowledge, it is necessary to introduce an individual at the appropriate position (key) to influence the discussion. To impact the conversation, this new individual must speak loudly —- both to counterbalance the voices of the others and to ensure that those positioned further away can hear them. This scenario can be further complicated by the presence of other nearby groups (representing distinct knowledge) that may be in close proximity to the original group. Consequently, the heightened vocal projection of the new participant has the potential to inadvertently influence the discussions occurring within these adjacent groups.
Figure 3: Left: The distribution of whitening similarity between two rephrases of a same subject. Right: A visualization for key representations of rephrases for ‘Slovenia’ in LLaMA2-7B.

### v2

@@ Abstract
Large language models (LLMs) struggle with maintaining accurate knowledge due to conflicting/outdated parametric memories. While locate-and-edit methods address this, their reliance on models’ internal representations leads to robustness failures in long-context reasoning and paraphrased queries. We identify a fundamental limitation of locate-and-edit methods: existing semantic keys (for memory localization) cannot simultaneously satisfy robustness (context-invariant activation) and specificity (precise knowledge discrimination). Through theoretical error-bound analysis, we establish formal criteria for effective editing. Our solution introduces Robust Edit Pathway (REP), a plug-and-play module that: (1) disentangles editing keys from native model representations; (2) dynamically adjusts keys via contrastive learning to achieve robustness-specificity balance. Extensive experiments across various editing methods (ROME/MEMIT/R-ROME/EMMET), existing LLMs (LLaMA2, QWen, Mistral), and datasets (CounterFact, ZsRE) show that REP improves success rate over robustness tests by up-to 66.4% while maintaining the success rate unaffected.  [^1] Our code can be found at https://github.com/ElliottYan/RobustKeyEdit.
Keys to Robust Edits: From Theoretical Insights to Practical Advances
Jianhao Yan1,2thanks:   Two authors contributed equally to this work.       Futing Wang1,211footnotemark: 1       Yun Luo1,2       Yafu Li4       Yue Zhang2,3thanks:   Corresponding author. 1Zhejiang University  2School of Engineering, Westlake University
3 Institute of Advanced Technology, Westlake Institute for Advanced Study
4 Shanghai AI Lab
elliottyan37@gmail.com
@@ 1 Introduction
Large language models (LLMs, Achiam et al. 2023; Touvron et al. 2023a, b) have revolutionized knowledge storage through their parametric memories, yet their reliance on static training data renders them prone to inaccuracies from conflicting or outdated information. While knowledge editing methods like ROME and MEMIT Meng et al. (2022a, b) attempt to address this by modifying specific model parameters, existing approaches are found to suffer from editing failures with robustness tests Ma et al. (2024c); Yang et al. (2024b). For example, editing "Slovenia belongs to Europe → Antarctica" frequently collapses when the subject is rephrased ("Republic of Slovenia"), embedded in long contexts, or attacked by shuffling subjects. The unreliability greatly limits the impact and application of model editing methods.
We uncover a fundamental flaw in their core mechanism: the intrinsic instability of the model’s internal representations when used as semantic keys for editing. Existing approaches assume these internal representations can reliably localize knowledge. Through formal analysis of key-value associative memory in MLP layers (Definitions 3.2 - 3.3) and empirical analyses, we prove that existing internal representations frequently violate the foundational conditions for reliable editing: (1) Key Sensitivity: Representations of the same fact diverge drastically under perturbations. Whitened similarity scores drop to near-random levels for shuffled subject tokens (e.g., "_ia Sloven" vs. "Sloven _ia") and for rephrased variants, breaching the robustness bound derived in Lemma 4.6; (2) Key Collisions: Semantically distinct entities exhibit unintended overlaps in the whitened space for unrelated pairs like "Michael Jordan" and "Kobe Bryant", Figure 4), contradicting the specificity requirement in Lemma 4.7.
@@
To resolve this, we propose Robust Edit Pathway (REP), a novel plug-and-play module that disentangles editing keys from native model representations. Inspired by our theoretical results, where effective knowledge insertion requires both centering around semantically equivalent surface forms of subjects while not affecting unrelated ones – REP introduces: (1) Disentangled Key Projection: A contrastively trained adapter aligns keys for target facts across perturbations, ensuring context-invariant activation through whitened similarity optimization (Eq. 6); (2) Dynamic Gate Mechanism: Token-level gating selectively activates edits, dynamically balancing robustness and specificity. Extensive evaluations across 4 editing methods (ROME/MEMIT/R-ROME/EMMET)(Meng et al., 2022a, b; Gupta et al., 2024a; Yoon et al., 2024), 3 LLMs (LLaMA2-7B, Mistral-7B, and Qwen-2-7B)(Touvron et al., 2023b; Jiang et al., 2023; Yang et al., 2024a), and two datasets (Meng et al., 2022a; De Cao et al., 2021) demonstrate REP’s superiority: (1) up-to 66.4% absolute gains on robustness tests, recovering at most 94% of the editing performance versus unperturbed inputs; (2) specificity preservation ($\Delta$Locality < 1.6) and minimal fluency degradation ($\Delta$Fluency <2.2); (3) effectiveness on both in-domain and out-of-domain robustness queries.
Our contributions are as follows:
  • Through theoretical error-bound analysis, we establish formal criteria for effective model editing and reveal fundamental limitations in using internal representations as editing keys.
  • Extensive experiments demonstrate existing semantic keys cannot simultaneously achieve robustness (context-invariant activation) and specificity (precise knowledge discrimination).
  • We propose Robust Edit Pathway (REP), a plug-and-play module that disentangles editing keys from native model representations and dynamically adjusts them via contrastive learning.
  • Experiments across various editing methods (ROME/MEMIT/R-ROME/EMMET), LLMs, and datasets show REP improves success rate over robustness tests by up-to 66.4% while maintaining editing performance.
@@ 2 Related Work
As large language models have grown in complexity and size, post-modification has become increasingly challenging due to their opaque mechanisms and vast parameter spaces (Mitchell et al., 2022; Zhong et al., 2023). This has led to heightened interest in knowledge editing, a technique for precise model modification. Knowledge editing are applied to various scenarios, such as editing for safety (Wang et al., 2024c), debias (Yan et al., 2024) and concepts (Wang et al., 2024d).
@@
Our work is in line with the locate-and-edit methods, which draw much attention as they potentially unveil how the knowledge are stored in an LLM. These approaches first identify relevant parameters before updating them to modify specific knowledge, including KnowledgeNeuron’s attribution-based neuron updating (Dai et al., 2021), ROME’s causal mediation analysis for MLP editing (Meng et al., 2022a), MEMIT’s multi-layer residual distribution (Meng et al., 2022b), PMET’s refined allocation strategy (Li et al., 2024), and WilKE’s dynamic layer selection (Hu et al., 2024b) to reduce potential negative effects. These methods all utilize inner representations as keys for key-value modeling. In contrast, we show that inner representations cannot meet the requirements of robust and specific edits, and we propose a robust edit pathway to mitigate this.
Challenges of Knowledge Editing.
@@
Despite the promise, various challenges persist in practical applications of model editing methods. Previous studies show that edits often degrade general language abilities (Gu et al., 2024; Ma et al., 2024b), damage the hidden space (Wang et al., 2024b), struggle to propagate to related facts (Hua et al., 2024), and are easily forgotten during sequential updates (Gupta et al., 2024b). Moreover, multi-hop reasoning can elicit old knowledge (Zhang et al., 2024), and models may collapse after few edits (Yang et al., 2024b; Brown et al., 2023).
@@
Further complications include cross-lingual inconsistencies (Wang et al., 2024a), knowledge conflicts (Li et al., 2023), and inadequate evaluation in realistic settings such as long-form generation (Rosati et al., 2024) and neighborhood knowledge (Ma et al., 2024a). These issues underscore the need for more sophisticated and comprehensive editing techniques. However, previous research largely remains focused on the outcomes of knowledge editing in various scenarios, lacking a deeper understanding of the underlying mechanisms of these methods and the true reasons behind their frequent failures. Our work presents both theoretical and empirical understanding regarding the reason for the robustness failures of locate-and-edit methods and proposes REP to enhance them.
Figure 2: Overview of REP. Left: Key concept visualization; Right: Architectural design of the adapter.
@@ 3 Knowledge Editing
In this section, we first formulate knowledge editing and review the locate-and-edit methods.
@@
Knowledge editing focuses on updating factual associations in language models. Following Meng et al. (2022a) and Meng et al. (2022b), we define a knowledge $\mathbf{f}$ as a triple $(h,r,t)$, where $h$ is the head entity, $r$ is the relation, and $t$ is the target entity (e.g., (USA, has president, Biden)). Given a knowledge triple: $(h,r,t)$, the goal is to modify the model’s knowledge by replacing the target entity $t$ with a new target $t_{*}=\text{Trump}$ (e.g., changing ‘Biden’ to ‘Trump’).
Autoregressive large language models (LLMs) can complete a natural-language sentence by leveraging implicit knowledge encoded within their parameters. Thus, a knowledge triple $(h,r,t)$ is considered stored in the LLM when the model can predict the target $t$ given a prompt that corresponds to $(h,r,\cdot)$. For instance, given a prompt ‘The president of USA is’, a model with the above knowledge would predict ‘Biden’.
Definition 3.1 (Knowledge Editing for LLMs).
Given a knowledge triple $(h,r,t)$ already stored in the language model $\mathcal{M}$ and a new knowledge $(h,r,t_{*})$, there exists a set of prompts $P=\{p\}$ corresponds to $(h,r,\cdot)$. The knowledge editing algorithm $\mathcal{A}$ aims to modify the model’s prediction on $P$ from $t$ to $t_{*}$. This task can be formally expressed as follows:
$\mathcal{M}^{\prime}=\mathcal{A}(\mathcal{M})\,,$
[…]
Architectural Foundations for locate-and-edit
The efficacy of locate-and-edit methods relies on identifying modular components in LLMs that encode factual knowledge. Transformer-based  Vaswani (2017) LLMs organize computation into layers containing two core submodules: self-attention (for contextual reasoning) and Multi-Layer Perceptrons (MLPs, for nonlinear feature transformations). A key insight from ROME (Meng et al., 2022a) establishes that factual associations localize to specific MLP layers—enabling precise edits.
Each MLP layer comprises two feed-forward operations: (1) an up-projection that expands hidden dimensions for fine-grained feature interactions, and (2) a down-projection that contracts dimensions to synthesize higher-level representations. ROME treats these MLPs as linear associative memories (Definition 3.2), leveraging causal mediation analysis to pinpoint layers where edits (e.g., substituting “Biden” → “Trump” in presidential facts) propagate correctly. By surgically modifying these layers, ROME updates targeted knowledge while preserving unrelated model capabilities, minimizing unintended side effects.
Definition 3.2 (MLP Layers as Associative Memories).
The down-projection weight matrix $W$ in the MLP layer can be interpreted as a linear associative memory system. Specifically:
  • Keys $K=[k_{1}|k_{2}|\cdots|k_{n}]\in\mathbb{R}^{D_{1}\times n}$ represent the intermediate representations of the prompt corresponding to $(h,r,\cdot)$ before down-projection.
  • Values $V=[v_{1}|v_{2}|\cdots|v_{n}]\in\mathbb{R}^{D_{2}\times n}$ represent the corresponding outputs after down-projection.
The weight matrix $W\in\mathbf{R}^{D_{2}\times D_{1}}$ approximately maps the keys to their associated values, satisfying $WK\approx V$.
Definition 3.2 (illustrated by Figure 2 left) enables the MLP layers to store and retrieve prompt-target associations. Then, ROME accomplishes knowledge editing by inserting a new key-value pair into the MLP layer, modifying $W$ to ${\hat{W}}$.
@@
$\text{minimize}_{\hat{W}}\ ||\hat{W}K-V||~{}\text{s.t.}\ \hat{W}k_{*}=v_{*}\,,$
by $\text{by setting}\ \hat{W}=W+\boldsymbol{\Lambda}(C^{-1}k_{*})^{T}$
@@
Intuitively, $||\hat{W}K-V||$ controls the shift from previously stored keys and values, and $\hat{W}k_{*}=v_{*}$ makes sure that the new knowledge is added into $\hat{W}$. To implement this solution, it is necessary to extract the key $k_{*}$ and calculate the value $v_{*}$.
@@
$v_{*}=\arg\min_{\mathbf{z}}\frac{1}{N}\sum_{j=1}^{N}\underbrace{-\log\mathbb{P}_{\mathcal{M}(m_{i}^{(l*)}:\mathbf{z})}[h^{\prime}|x_{j}+p]}_{\text{(a) Maximizing }h^{\prime}\text{ probability}}$
$+\underbrace{D_{\text{KL}}\left(\mathbb{P}_{\mathcal{M}(m_{i}^{(l*)}:\mathbf{z})}[\cdot|p^{\prime}]||\mathbb{P}_{\mathcal{M}}[\cdot|p^{\prime}]\right)}_{\text{(b) Controlling essence drift}}.$
where $p^{\prime}$ is ‘subject is a’.
@@ 4 Theoretical Results of Key-Value Associative Memory
The idea of keys and values in associative memory (as shown in Definition 3.2) is analogous to the key-value databases in modern computer systems. What makes the difference here is that down-projection FFNs implement a fuzzy retrieval mechanism, whereas modern key-value databases generally require the keys to be unique.
@@
Given $K\in\mathbb{R}^{D_{1}\times n}$ and $V\in\mathbb{R}^{D_{2}\times n}$ as defined in Definition 3.2 that are already stored in the feed-forward layer $W\in\mathbb{R}^{D_{2}\times D_{1}}$, assume $n\gg D_{1}$ and $K$ has the rank of $D_{1}$. When a new query $k_{*}$ comes, its corresponding value can be represented as the weighted sum of existing values, $v_{*}=\sum^{N}_{{i=0}}\alpha_{i}v_{i}$ and $\alpha=K^{T}(KK^{T})^{-1}k_{*}$ can be solved by the Moore-Penrose pseudoinverse.
Lemma 4.1 demonstrates that the retrieved memory of a new test query can be considered as the linear combination of previously stored memory, which leads to the following direct corollary.
@@
In locate-and-edit algorithms, new knowledge is injected into the memory as a key-value pair $(k_{*},v_{*})$. Consider a set of existing key-value pairs $(k_{i},v_{i})$ where $k_{i}\in\mathcal{K}_{s},v_{i}\in V_{s}$ that represent the same knowledge as $(k_{*},v_{*})$ (e.g., paraphrases). Suppose the injection is lossless [^2] In a real-world scenario, the edit cannot be lossless. Here, for a clear intuition, the above lemma is presented in an ideal way as the editing process will change the value of previously stored key-value pairs. We show that even considering the lossless scenario, the current LLMs cannot satisfy robustness and specificity requirements. and that $\mathcal{K}_{s}$ has full row rank, querying with any $k_{i}\in\mathcal{K}_{s}$ would retrieve a value $v=\sum_{i}\alpha_{i}v_{i}+\alpha_{*}v_{*}$.
@@
This corollary reveals that knowledge editing operates as an additive mechanism rather than a replacement one. Instead, it leaves the previously stored knowledge intact and counters them with a newly added value $v_{*}$.
@@
Lemma 4.4 suggests that an edited value should be first similar to the vector pointing from $t$ to $t_{*}$ after a projection with $W_{V}$. Then, the edited values $\Delta v$ should be sufficiently large to ensure the success of the edit.
Our assumption here simplifies the connection between the edited layer and the prediction layer, as in real-world scenarios, the edit layer might pass through subsequent layers and undergo multiple attention operations before finally connecting to the prediction layer. However, the path we’re considering (i.e., from the edit layer to the prediction layer via an attention layer) is arguably the most direct route. We contend that this direct path is crucial and warrants particular attention, and this simplification allows us to focus on the most immediate and potentially significant impact of edits.
@@
Robust editing requires consistency across semantically equivalent inputs: when editing knowledge with a new pair $(k_{*},v_{*})$, the edit should propagate to all semantically equivalent representations in the memory. For an edit to be considered a robust edit, querying with any $k_{s}\in\mathcal{K}_{s}$ should reliably retrieve the new knowledge $(h,r,t_{*})$. This can be expressed as the following constraint:
$(w_{t_{*}}-w_{t})^{T}W_{V}(k_{s}^{T}C^{-1}k_{*})\cdot v_{*}^{T}>\epsilon_{1}+\epsilon_{2},\forall k_{s}\in\mathcal{K}_{s}.$
When we look into the Lemma 4.6, $\beta_{s,*}=k_{s}^{T}C^{-1}k_{*}$ can be seen as a similarity measure on a projected space, namely whiten similarity. This lemma implies that (1) $v_{*}$ is decided by $\min_{k_{s^{\prime}}\in\mathcal{K}_{s}}(k_{s^{\prime}}C^{-1}k_{*}$), that is, $k_{*}$ should be near all $k_{s}\in\mathcal{K}_{s}$. If not, $v_{*}$ needs to be of large magnitude to counter the difference. Such large-magnitude updates can destabilize the model’s learned representations and potentially degrade its overall performance; (2) $v_{*}$ needs not only to be aligned with the direction $(w_{t_{*}}-w_{t})$, but also has a sufficiently large magnitude to ensure editing success.
@@
If the newly added knowledge triplet $(h,r,t_{*})$ would not be retrieved for any $k_{o}\notin\mathcal{K}_{s}$, it requires the following inequality to be satisfied:
$(w_{t_{n}}-w_{t_{*}})^{T}W_{V}(k_{o}^{T}C^{-1}k_{*})\cdot v_{*}^{T}<\epsilon_{3},$
$\forall k_{o}\notin\mathcal{K}_{s}~{}and~{}\forall w\in W,$
where $t_{n}$ is the original target retrieved by $k_{o}$ and $\epsilon_{3}$ denotes the logit difference between $t_{n}$ and $t_{*}$.
One simple solution for this lemma is $k^{T}C^{-1}k_{*}=0$, which describes no superposition, as discussed in one of the concurrent work (Hu et al., 2024a). However, as superposition generally exists among existing LLMs, we discuss more general cases here.
Lemma 4.8 (Whitened Similarity Bounds).
For a successful edit to achieve both robustness and specificity, the whitened similarities must satisfy:
  1. 1.
    Lower bound for semantically equivalent keys:
    $\beta_{s,*}=k_{s}^{T}C^{-1}k^{*}\geq\beta_{\min},\quad\forall k_{s}\in\mathcal{K}_{s}$ (1)
    where $\beta_{\min}=\frac{\epsilon_{1}+\epsilon_{2}}{||(w_{t^{*}}-w_{t})^{T}W_{V}||\cdot||v^{*}||}$
  2. 2.
    Upper bound for unrelated keys:
    $|\beta_{o,*}|=|k_{o}^{T}C^{-1}k^{*}|\leq\beta_{\max},\quad\forall k_{o}\notin\mathcal{K}_{s}$ (2)
    where $\beta_{\max}=\frac{\epsilon_{3}}{\max_{w\in\mathcal{W}}||w-w_{t^{*}}||\cdot||W_{V}||\cdot||v^{*}||}$
Detailed proof of all lemmas can be found in Appendix A.
Remark 4.9.
Lemma 4.9 suggests that when adding new knowledge, a new key must be introduced at an appropriate position. This new key must be placed carefully, as its position can affect both its intended target and potentially interfere with nearby keys.
Figure 3: The distribution of normalized whitening similarity between different kinds of keys and original keys.